OpenStudy (anonymous):

2tan(inv)a + 2 tan(inv)b = 2tan(inv)x .. (A) (1-b)/1-ab (B) (1+ab)/a-b (C) ab-1/a+b (D) a-b/1+ab

OpenStudy (anonymous):

u want the value of x?

OpenStudy (anonymous):

yes brother

OpenStudy (anonymous):

simply apply the formula

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

i got it thanx

OpenStudy (anonymous):

what i am getting is: (a+b)/(1-ab)

OpenStudy (anonymous):

@ishaan

OpenStudy (anonymous):

i am sorry , there is actually a minus sign not plus sign

OpenStudy (anonymous):

answer is d

OpenStudy (anonymous):

oh.. then D is correct.. it was making me confused..

OpenStudy (apoorvk):

check the questions and the options once again correctly. the "signs'. should be tan.invA "-" tan.invB. then the answer could be D

OpenStudy (apoorvk):

now i read the comments.

OpenStudy (anonymous):

first divide both sides by 2 to have \[\tan^{-1}a+\tan^{-1}b=\tan^{-1}x\] then get the tan of both sides to \[\tan(\tan^{-1}a+\tan^{-1}b)=\tan(\tan^{-1}x)\] \[\tan(\tan^{-1}a+\tan^{-1}b)=x\] then using the trigonometric angle sum identities: \[\tan(a+b)=(\tan a+\tan b)/(1-tana tanb)\] we find x, so the LHS now becomes \[(\tan(\tan^{-1}a)+\tan(\tan^{-1}b))/(1-\tan(\tan^{-1}a)\tan(\tan^{-1}b))\] \[x=(a+b)/(1-ab))\] so I have to agree with @apoorvk this will be the answer IF: 2tan(inv)a + 2 tan(inv)b = 2tan(inv)x D will be the answer IF:2tan(inv)a - 2 tan(inv)b = 2tan(inv)x